# HMSViT: A Hierarchical Masked Self-Supervised Vision Transformer for Corneal Nerve Segmentation and Diabetic Neuropathy Diagnosis

### v2

@@ 6 Discussion
This approach addressed key challenges in medical imaging, where limited labelled data often restricts deep learning applications. Unlike previous models that relied on large annotated datasets, our SSL approach enabled efficient learning from limited labelled data. This was especially valuable for high-dimensional corneal confocal microscopy (CCM) images, which were often scarce. Comparisons with state-of-the-art methods highlighted HMSViT’s superior performance over models using only hierarchical vision transformers or convolutional neural networks. By leveraging multi-scale feature extraction, our model captured both fine-grained local details and broader structural patterns, enhancing segmentation accuracy and improving diagnostic reliability. The use of multi-scale feature extraction enabled the proposed model to discern both fine-grained local features and broader structural patterns, improving segmentation accuracy and facilitating the extraction of clinically relevant parameters, such as corneal nerve fibre length (CNFL) and branch density.
Figure 5: Multi scale feature extraction from the proposed HMSViT.
@@
One of the key innovations of the HMSViT model is its ability to capture hierarchical multi-scale features, effectively combining local and global contextual information. This capability is crucial in medical imaging tasks where both fine-grained details and broader structural patterns contribute to accurate diagnosis. As illustrated in Figure 5, the hierarchical feature extraction approach enabled the model to enhance segmentation accuracy, as evidenced by its superior performance in mean Intersection over Union (mIoU) compared to existing models. The hierarchical nature of the HMSViT allowed for better retention of nerve fibre structure and branching patterns, ensuring more precise morphological assessments. Another innovation in HMSViT is its streamlined design featuring a pooling-based token mixer, which minimizes architectural complexity while maintaining robust performance. This simplicity is key to the model’s computational efficiency. Notably, HMSViT achieved an mIoU of 61.34% using only 68M parameters, outperforming more complex architectures that required significantly higher computational resources. Such efficiency is critical for deployment in clinical environments where processing speed and resource constraints are paramount. A significant challenge in medical imaging is the scarcity of labelled datasets, which limits the performance of supervised deep learning models. The proposed HMSViT mitigates this limitation by leveraging self-supervised learning, which enables the model to learn meaningful representations from large amounts of unlabelled data. The ablation study results confirmed that the addition of self-supervised learning significantly enhanced both segmentation and classification performance. Notably, the model trained with SSL and the hierarchical structure achieved the highest classification accuracy of 70.40% and an mIoU of 61.34%. These improvements underscore the importance of SSL in enabling efficient feature learning, reducing reliance on expensive manual annotations.
@@
A direct comparison of the proposed model with other state-of-the-art methods highlights its superior performance. While previous hierarchical vision transformers such as HiViT and Fast-iTPN achieved competitive results, the HMSViT surpassed these models in segmentation performance due to its optimized pyramid structure and modified Masked Autoencoder (MAE) approach. The ability to outperform models that rely solely on hierarchical architectures or traditional convolutional networks demonstrates the effectiveness of integrating SSL with multi-scale feature extraction. The proposed approach not only improves segmentation quality but also enhances diagnostic accuracy, reinforcing its potential for real-world clinical applications.
@@ 7 Conclusion
In this work, we presented HMSViT—a Hierarchical Multi-Scale Vision Transformer that integrates a novel block masked self-supervised learning strategy with a streamlined hierarchical architecture. By combining a pooling-based hierarchical design with a multi-scale decoder, HMSViT successfully captures both fine-grained local details and broader global contexts from corneal confocal microscopy images. Our extensive experiments on clinical datasets demonstrated that HMSViT achieves state-of-the-art performance, with a mean Intersection over Union (mIoU) of 61.34% for corneal nerve segmentation and a diagnostic accuracy of 70.40% for diabetic peripheral neuropathy classification. Ablation studies confirmed that the improvements are driven by both the block masked self-supervised pretraining and the hierarchical feature extraction, allowing HMSViT to outperform competitive models like the Swin Transformer and HiViT while utilizing fewer parameters.
@@
[…]

### v3

@@ 6 Discussion
This approach addresses key challenges in medical imaging, where limited labelled data often constrains the effectiveness of deep learning methods. Unlike previous models that rely heavily on large annotated datasets, our self-supervised learning (SSL) strategy enables efficient representation learning from limited labelled data. This is particularly valuable for high-dimensional corneal confocal microscopy (CCM) images, which are often scarce and expensive to annotate. Comparisons with state-of-the-art methods highlight HMSViT’s superior performance over models relying solely on hierarchical ViTs or CNNs. The proposed model’s multi-scale feature extraction enables the simultaneous capture of fine-grained local details and broader structural context, which is critical for accurate segmentation. This capability not only improves diagnostic reliability but also facilitates the extraction of clinically relevant parameters, such as corneal nerve fibre length (CNFL) and branch density.
Figure 7: Multi-scale feature representations extracted from a Vision Transformer model applied to corneal confocal microscopy images. The first column shows the original input images. The second, third, and fourth columns illustrate feature maps at different spatial resolutions (256×256, 128×128, and 64×64), capturing hierarchical representations from fine to coarse levels. Fine features highlight detailed nerve structures, while coarser features capture more abstract patterns.
@@
One of the key innovations of the HMSViT model is its ability to capture hierarchical multi-scale features, effectively combining local and global contextual information. This capability is crucial in medical imaging tasks where both fine-grained details and broader structural patterns contribute to accurate diagnosis. As illustrated in Figure 7, the hierarchical feature extraction approach enables the model to enhance segmentation accuracy, as evidenced by its superior performance in mean Intersection over Union (mIoU) compared to existing models. The hierarchical nature of the HMSViT allows for better retention of nerve fibre structure and branching patterns, ensuring more precise morphological assessments. Another innovation in HMSViT is its streamlined design featuring a pooling-based token mixer, which minimizes architectural complexity while maintaining robust performance. This simplicity is key to the model’s computational efficiency. Notably, HMSViT achieved an mIoU of 61.34% using only 68M parameters, outperforming more complex architectures that required significantly higher computational resources. Such efficiency is critical for deployment in clinical environments where processing speed and resource constraints are paramount.
A significant challenge in medical imaging is the scarcity of labelled datasets, which limits the performance of supervised deep learning models. The proposed HMSViT mitigates this limitation by leveraging self-supervised learning, which enables the model to learn meaningful representations from large amounts of unlabelled data. The ablation study results confirmed that the addition of self-supervised learning significantly enhanced both segmentation and classification performance. Notably, the model trained with SSL and the hierarchical structure achieved the highest classification accuracy of 70.40% and an mIoU of 61.34%. These improvements underscore the importance of SSL in enabling efficient feature learning, reducing reliance on expensive manual annotations.
@@
A direct comparison of the proposed model with other state-of-the-art methods highlights its superior performance. While previous hierarchical vision transformers such as HiViT and Fast-iTPN achieved competitive results, the HMSViT surpassed these models in segmentation performance due to its optimised pyramid structure and modified Masked Autoencoder (MAE) approach. The ability to outperform models that rely solely on hierarchical architectures or traditional convolutional networks demonstrates the effectiveness of integrating SSL with multi-scale feature extraction. The proposed approach not only improves segmentation quality but also enhances diagnostic accuracy, reinforcing its potential for real-world clinical applications.
@@ 7 Conclusion
In this work, we presented HMSViT, a Hierarchical Multi-Scale Vision Transformer that integrates a novel block-masked self-supervised learning strategy with a streamlined hierarchical architecture. By combining a pooling-based hierarchical design with a multi-scale decoder, HMSViT successfully captures both fine-grained local details and broader global contexts from corneal confocal microscopy images. Our extensive experiments on clinical datasets demonstrated that HMSViT achieves state-of-the-art performance. Notably, it attained a patient-level diagnostic accuracy of 85.60% and a segmentation mIoU of 61.34%, outperforming leading models like the Swin Transformer and HiViT. This superior accuracy was achieved while using significantly fewer parameters—for instance, approximately 41% fewer than the Swin Transformer. Ablation studies further confirmed that the synergy between our hierarchical design and block-masked pretraining is critical to its success, driving the substantial improvements in both segmentation and classification tasks.
@@
CRediT authorship contribution statement
Xin Zhang: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Visualization, Writing – original draft. Liangxiu Han: Conceptualization, Methodology, Supervision, Project administration, Funding acquisition, Writing – review & editing. Yue Shi: Software, Validation, Writing – review & editing. Yalin Zheng: Data curation, Resources, Project administration, Funding acquisition, Writing – review & editing. Uazman Alam: Data curation, Resources, Writing – review & editing. Maryam Ferdousi: Data curation, Resources, Writing – review & editing. Rayaz Malik: Conceptualization, Data curation, Resources, Writing – review & editing.
Acknowledgments
This work was supported by the Engineering and Physical Sciences Research Council (EPSRC) through grants EP/X013707/1 and EP/X01441X/1.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability
The datasets generated and analysed during the current study are not publicly available due to patient privacy restrictions but are available from the corresponding author on reasonable request. All studies contributing to the clinical dataset received ethical approval from the NRES Committee North West - Greater Manchester committees, including the NIH study (Ophthalmic Markers of Diabetic Neuropathy, REC number: 09/H1006/38), the LANDMARK study (Surrogate markers for diabetic neuropathy, REC number: 08/H1004/1), and the PROPANE study (Probing the Role of Sodium Channels in Painful Neuropathies, REC number: 14/NW/0093). Written informed consent was obtained from all participants.
Code Availability
The source code for this project is publicly available at https://gitlab.com/han-research/hmsvit.